Summary EVE shift Tues. 8/24/99

Richard Ford (rickford@fnal.gov)
Tue, 24 Aug 1999 23:58:57 -0500 (CDT)

Shift Summary for EVENING shift, Tuesday Aug 24, 1999

On Shift: Rick Kessler and Rick Ford

Hours of beam delivered at all intensities: 7.5 hrs
Hours of beam delivered at nominal intensity: 7.5
Hours of E832 data (special runs not included): 5.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
28 minutes : Changing tapes, stop-start new run
1 hr 30 minutes: ETOT LV supply swap

========================================================================
Runs ended this shift:
Run 14319 BEAM_832 - 14.5k vac 2pi0 QK221-223, QKB112 taken to FCC
14320 Beam_832 - 4.3k vac 2pi0
14322 Beam_832 - in progress
========================================================================
Accesses during shift: Replace ETOT LV supply

========================================================================
Overview of shift: Fairly smooth until 21:23 when we received an alarm on
ETOT LV5 reulting in an access and swapping of Power
supply.
Smooth sailing before and after

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14319 15:30 Inherit BEAM_832 run from day shift

7,8 15:40 Beam returns; notice NM3S is not ramping. MCR sets it
to ramp

168 19:15 Beam goes away. Stop run 14319. Change tapes
STOP button hangs, triggers not stopped, waited for
CsI gain process to finish, drt_fresh_start

19:20 Beam returns

14320 01 19:38 Start run 14320, HCC illumination, upper right
quadrant blank. Re-load ETOT.

50 20:43 Beam disappears. Beam has been sporadic on Booster
long 7 rad trips for the last 45 min - 1hr. Replacement
of interlocked detector underway - est 15 min.

72 21:17 Beam returns.

78 21:23 ETOT LV 5 trip. 2 resets failed. Experts enroute.

14321 22:01 init COSMICS after modifying destination so that
HCC is read out for COSMICS. Should see hot board
if LV is still dead.
Jason & Sydney are on access replacing LV supply

22:45 Sydney & Jason return from access, request beam

14322 22:55 Start run
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift